OpenStudy (mysticfury):

Change this to slope - intercept form :) medal and fan ! 2x − 5y = −15

OpenStudy (astrophysics):

Slope intercept form is y = mx+b, so solve for y.

OpenStudy (mysticfury):

well to do that i would need to know how im not great at solving for y in this :/

rebeccaxhawaii (rebeccaxhawaii):

move x to the other side

OpenStudy (astrophysics):

Try subtracting 2x from both sides, what happens?

OpenStudy (mysticfury):

5y = -15 - 2x ?

OpenStudy (mysticfury):

or is it + instead of -?

OpenStudy (lexiquesse):

id put the -2x before -15

OpenStudy (freeboy15):

+15 to both sides

OpenStudy (lexiquesse):

its also a -5y

OpenStudy (astrophysics):

Here, \[\color{red}{-2x}+2x-5y=-15 \color{red}{-2x}\] notice what I do to one side I also do it to the other, this really shows that the equation is still the same, but we're just changing the location of the variables/ numbers.

OpenStudy (astrophysics):

So it will get cancelled on the left side -2x+2x = 0 right?

OpenStudy (mysticfury):

yes i already knew that @Astrophysics

OpenStudy (nincompoop):

on both sides, +15, +5y then solve for y

OpenStudy (astrophysics):

Now we have \[-5y=-15-2x\] what can we do?

OpenStudy (mysticfury):

minus 15 ? @Astrophysics

OpenStudy (nincompoop):

it does not matter whether the y is on the right side or left side. 3 = y and y = 3 are the same

OpenStudy (lexiquesse):

so you have -5y=-2x-15. first multiply -2x-15 by -1 -(-2x-15)

OpenStudy (lexiquesse):

then divide that by 5

OpenStudy (astrophysics):

Not quite Mystic we want to solve for y, so we can leave the right side alone, solving for y means we want y by itself. So notice that the -5 is "attached" to the y there, which means it's being multiplied by the y, so what it means is \[-5 \times y = -15-2x\] that's what is going on, but it looks convoluted so we write -5y. Now think to yourself what we can do here, your best bet would be to use the opposite operation of what already exists, so in this case we can use division. \[\frac{ -5y }{ \color{red}{-5} }=\frac{ -15-2x }{ \color{red}{-5} }\]

OpenStudy (mysticfury):

okay so would we take the -15 divided by -5? or take -15 - 2 then divide? @Astrophysics

OpenStudy (astrophysics):

We divide the whole right side by -5

OpenStudy (astrophysics):

\[y=\frac{ \left( -15-2x \right) }{ -5 }\] maybe this is more clear, you can simplify it more if you like

OpenStudy (mysticfury):

umm okay i just 3.4x ;/

OpenStudy (mysticfury):

how do you divide that?

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

it is actually easier to work with when your slope is in fractional form

OpenStudy (astrophysics):

\[y=\frac{ \left( -15-2x \right) }{ -5 } \implies y = \frac{ -15 }{ -5 }-\frac{ 2x }{ -5 }\] you can break it up as such, you will have it in the form y = mx+b

OpenStudy (astrophysics):

Can you simplify that?

OpenStudy (mysticfury):

3 + -2.5x or -0.4x?

OpenStudy (astrophysics):

\[y=3+\frac{ 2 }{ 5 }x ~~~ or ~~~ y= \frac{ 2 }{ 5 }x+3\]

OpenStudy (mysticfury):

ah gotcha, you put it in decimal form, i have a question though, would 2/5 be a positive or negative slope?

OpenStudy (astrophysics):

It's positive, if it was negative there would be a negative sign indicating so, remember slope is rise/run up 2 and 5 right

OpenStudy (mysticfury):

okay i wasnt sure ,since its less then 1, thanks so much @Astrophysics !

OpenStudy (astrophysics):

Yw